OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, -6) and is perpendicular to y = 3x - 6.

OpenStudy (lalaly):

since its perpindicular that means it has the slope -1/3 so y--6=-1/3 (x-6) y+6=-1/3 x +2 y=-1/3 x-4

OpenStudy (anonymous):

Thankkyuu

OpenStudy (lalaly):

yur welcome